OpenStudy (anonymous):

What is the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3)? How would I find it? I looked it up and it didn't make sense. :( Medal and fan!!

OpenStudy (michele_laino):

please you have to use the subsequent formula: \[y-(-3)=7(x-1)\] where I used the slope of the line in the text of your problem, because the request is another line which is parallel to the first line, namely y=7x+2

OpenStudy (michele_laino):

please complete the calculations!

OpenStudy (anonymous):

y=7x+b: -3=7*1+b b=-10 y=7x-10